2 4.2 m A 55 kg person stands on a uniform 17 kg, 6.0 m long ladder resting against a frictionless wall EXAMPLE NOTES IMAGES DISCUSS UNITS STATS HELP Part Description Answer Chk History What is the magnitude of the force of the wall on ladder? (include units with answer) 6.67 pts. 100% 2% try penalty Hints: 2, # tries: 0 Show Details Format Check What is the magnitude of the normal force of the ground on ladder? (include units with answer) 6.67 pts. 100% 12% try penalty Hints: 1,0 # tries: 0 Show Details Format Check What is the minimum coefficient of friction so the ladder does not slip? 6.67 pts. 100% 2% try penalty Hints: 0,0 # tries: 0 Show Details What is the minimum coefficient of friction so the ladder does not slip for any location of the person on the ladder? 6.67 pts. 100% 2% try penalty Hints: 2, # tries: 0 Show Details
Answer #1

given

masses are 55 kg and 17 kg

6 m long ladder

and 4.2 m , angle 75o

a )

the magnitude of the force of the wall on ladder is

17 x 3 x cos75 + 55 x 4.2 x cos75 = Rw x 6 x sin75

Rw = 17 x 3 x cos75 + 55 x 4.2 x cos75 / 6 x sin75

Rw = 12.593 kg

b )

the magnitude of the normal force of the ground on ladder is RH = 17 + 55 = 72 kg

c )

the minimum coefficient friction is

\mu x 72 \geq 12.593

\mu\geq 12.593 / 72

\mu\geq 0.1749

or

\mu\geq 0.175

d )

the minimum coefficient friction is

17 x 3 x cos75 + 55 x 6 x cos 75 = Rw x 6 x sin75

Rw = 17 x 3 x cos75 + 55 x 6 x cos 75 / 6 x sin75

Rw = 17.015 kg

\mu = Rw / 72

= 17.015 / 72

\mu = 0.2363